# TESSERACT Collaboration Low Energy Backgrounds and Excess Noise in a Two-Channel Low-Threshold Calorimeter

### v1

@@ Abstract
We describe observations of low energy excess (LEE) events (background events observed in all light dark matter direct detection calorimeters) and noise in a two-channel silicon athermal phonon detector with 375 meV baseline energy resolution. We measure two distinct LEE populations: “shared” multichannel events with a pulse shape consistent with athermal phonon events, and sub-eV events which couple nearly exclusively to a single channel with a significantly faster pulse shape. These “singles” are consistent with events occurring within the aluminum athermal phonon collection fins. Similarly, our measured detector noise is higher than the theoretical expectation. Measured noise can be split into an uncorrelated component, consistent with shot noise from small energy depositions within the athermal phonon sensor itself, and a correlated component, consistent with shot noise from energy depositions within the silicon crystal’s phonon system.
I Introduction
@@
The search for dark matter (DM) has expanded to lower mass candidates, including sub-GeV “light mass” DM Kuflik et al. (2016, 2017); Hochberg et al. (2014, 2015); Hall et al. (2010). Direct detection of light mass DM scattering off nuclei, electrons, or crystal lattices requires extremely low energy thresholds, given the low kinetic energy carried by the DM particles. Cryogenic calorimeters are well suited to attaining such low thresholds, and have recently set limits on sub-GeV DM-nucleon interaction cross sections Abdelhameed et al. (2019); Alkhatib et al. (2021); Angloher et al. (2023). These calorimeters typically read out athermal phonons from a crystal target using Transition Edge Sensors (TES) Irwin and Hilton (2005) connected to aluminum athermal phonon collection fins (forming structures known as Quasiparticle-trap-assisted Electrothermal-feedback TESs: QETs Irwin et al. (1995)).
@@
Calorimetric DM direct detection experiments and other low threshold calorimeters have observed an excess of events below several hundred eV, with a rate that rises dramatically at low energies Adari et al. (2022); Alkhatib et al. (2021); Abdelhameed et al. (2019). The rate of these low energy excess (LEE) events decreases with time, and can be regenerated by warming up the detector Angloher et al. (2022). Additionally, the LEE rate varies only weakly with detector material or mass Angloher et al. (2022), and appears similarly in detectors run above and below ground.
@@
The decrease in LEE rate with time suggests a relaxation mediated process. Mechanical stress relaxation in the detector holding has been shown to create LEE-like events Astrom et al. (2006); Anthony-Petersen et al. (2022); however, a LEE population remains even in detectors held in low stress configurations Anthony-Petersen et al. (2022), implying additional relaxation processes are necessary to explain observations.
@@
Stress created by the thermal contraction of sensor films relative to thick detector substrates has been proposed as another LEE source Anthony-Petersen et al. (2022); Romani (2024). This stress would be present in all calorimeters observing the LEE, largely independent of the detector material or size.
@@
If relaxation within films is responsible for the LEE, we expect some partitioning of energy between local deposition in the film’s electronic system and phonon energy that leaks into the substrate. For example, athermal phonon bursts within the aluminum QET should break Cooper pairs and locally deposit energy Kozorezov et al. (2000); Martinis (2021). A detector instrumented with multiple individually-read-out athermal phonon sensors would measure this localized energy deposition and therefore be able to discriminate LEE events (with significant local energy absorption) from DM interactions in the bulk substrate (that approximately uniformly excite all phonon sensors). This detector architecture has previously been shown by CRESSTAngloher et al. (2024) and contemporaneously by our group Romani (2023) to have great potential. However, limited control of systematics and understanding of pulse shapes for locally absorbed events constrained our ability to draw strong conclusions about sources of LEE. These limitations have now been remedied.
II Detector Design and Data Collection
@@
To test this LEE discrimination concept, a 1 cm-square, 1 mm-thick silicon crystal was instrumented with two channels of 25 tungsten TESs ( 48 mK Tc) connected to aluminum athermal phonon collection fins (QETs Irwin et al. (1995)), covering 1.38 of the device’s surface. QETs were electrically connected by partially overlapping their aluminum fins. Unfortunately, during manufacturing, a fraction of the TESs were partially etched away, leading to some performance degradation (higher normal resistance and worse phonon collection efficiency than previous similar devices, suggesting some of the TESs within the channel are damaged). While performance in both channels was still acceptable, we focus on the left channel, which was less negatively impacted. As in Ref. Anthony-Petersen et al. (2022), this detector was suspended from wire bonds to minimize LEE-type backgrounds associated with detector holding. It was housed inside multiple layers of electromagnetic interference (EMI) and infrared (IR) shielding at the base stage of a dilution refrigerator, and was read out using SQUID array amplifiers.
@@
The detector was calibrated with optical photons to characterize its response to events of a known energy (see Figure 1). Pulses of photons from a 405 nm (3.061 eV) room temperature laser were transmitted to the device using a single mode optical fiber terminated with a diffuser, which dispersed photons across the entire instrumented side of the detector. The photon pulses were fast ( 100 ns) compared to the electrical ( 10 s) and electrothermal ( 100 s) response times of the TESs. On average photons hit the detector per pulse. Immediately following the calibration, background data were acquired.
@@
Both datasets were recorded continuously and triggered offline. For the calibration dataset, we triggered on a recorded logic signal which pulses when the laser fires. In the background dataset, events were triggered on the sum of the two detector channels using an optimum filter energy estimator Golwala (2000).
@@
Standard quality cuts were applied to the triggered data to ensure that the detector was operating stably, and to reject periods of high environmental noise or abnormal device performance. These cuts were designed for high passage of randomly triggered events (79.5 %) and for similar passage of high energy events to minimize selection biases. The height of each event was fit with an optimal filter assuming a calibrated phonon pulse shape (see below), and was converted into an energy by applying a factor derived from the channels responsivity Romani et al. (2024) modeled from the measured complex impedance Lindeman et al. (2011); Fink, Watkins et al. (2020).
III Calibration and Background
@@
In the calibration and background datasets (see Figs. 1 and 2), “shared” events couple roughly equally to both channels (diagonal band, left panels). These events feature a relatively slowly rising pulse (see top right panels) which can be well-modeled as the sum of two exponential phonon pulses convolved with the detector responsivity (). Calibration and background events are identically shaped.
@@
We associate these events with bursts of athermal phonons from the substrate which couple roughly equally to both channels. In the background dataset at low energies, we associate these “shared events” with non-sensor film LEE relaxation sources due to the lack of significant localized energy absorption within the channel. At high energies (off plot in Fig. 2), the saturated event rate is roughly consistent with the expected rate of high energy events from environmental radioactivity and cosmic rays.
@@
In the calibration, these events are caused by photons absorbed in the substrate, creating quantized (0, 1, 2… photons absorbed) bursts of athermal phonons. We combine the response in both channels using inverse variance weighting, constructing a phonon energy estimator, and plot a combined calibration histogram (see Fig. 1, inset). From this histogram, we measure a world leading baseline phonon energy resolution of meV.
@@
Another class of events exhibits a nearly maximally asymmetric channel response (vertical and horizontal bands in Figs. 1 and 2), which we call “single” events due to their strong coupling to a single channel. In the calibration dataset, we attribute this response to events in which one photon hits an aluminum phonon collection fin, i.e. a “direct hit” event. Additional photons may be absorbed in the detector crystal, creating a superposition of “direct hit” and “substrate” events and forming the structure of black bands in Fig. 1. Of the roughly 4.1 single photon events seen in the substrate, we expect 1 will hit a fin in a given channel, with 10 of these photons being absorbedFlaschmann et al. (2023). 239 events fall into the purple box in Fig. 1 (which contains roughly half of the left channel singles), in general agreement with the expected number of events. We attribute the spread in reconstructed energy to instrumental effects (e.g. position dependence within the aluminum, saturation from partial TESs etching). Given only 0.1 of the QET area is exposed tungsten, we do not expect to observe a significant number of tungsten direct hits.
@@
Pulse shapes for single events are shown in the bottom right panels of Figs. 1 and 2. The fast rise compared to substrate events (though somewhat slower than the modeled TES response to Dirac-delta energy impulses) indicates that the substrate phonon collection dynamics are bypassed. We attribute the slow fall of these events to saturation effects. Quasiparticles created by localized photon absorption would be expected to propagate to only a few of the 25 QETs in a channel, saturating these few TESs at energies significantly below the 6 eV entire channel saturation energy.
@@
In addition to the large response in the primary channel (left channel in Fig. 3), we see a small response in the opposite (right) channel, carrying a few percent of the main channel’s energy. The opposite channel responses for background and calibration events are indistinguishable, and consistent with the calibrated athermal phonon response. We attribute this signal to athermal phonons which leak out of the aluminum fin during downconversion following a singles event. Notably, without rescaling the background and calibration events, the responses are roughly equal in height, suggesting that lower energy background events are more efficient in producing athermal phonons than calibration events, and hinting that background singles may originate deeper within the aluminum fin.
@@
Bursts of high frequency (MHz-GHz) EMI Romani et al. (2024) cannot be the primary source of these background singles, as individual EMI photons would not be sufficiently energetic to leak above the aluminum gap ( 90GHz) phonons into the substrate during the downconversion process. We also reject relaxation or other events occurring in the tungsten TESs as the source of background singles, as tungsten relaxation would be expected to produce a far larger pulse of athermal phonons in the substrate due to thinness of the tungsten film ( nm)Romani et al. (2024).
@@
Alternatively, the relaxation of thermally stressed aluminum films could create singles backgrounds. Ref. Romani (2024) describes a model where relaxing dislocations in an aluminum film impact the aluminum-substrate interface, injecting bursts of athermal phonons into the substrate while depositing minimal energy in the aluminum. Clearly, observed singles are in the opposite limit: they couple almost exclusively to the aluminum film. Modifying the model in Ref. Romani (2024) to include e.g. damping of dislocations in the bulk film through the emission of above-gap phonons or phonon-emitting interactions with intra-film grain boundaries might better explain singles events.
Background spectra for single and shared events are plotted in Fig. 4. Shared events were triggered on the sum of the two channels using a phonon template, while singles were triggered in the left channel using an averaged singles template. Since a given single or shared event could trigger both single and shared triggers, a statistic considering the pulse shape and amplitude in both channels was used to discriminate event types and avoid double counting. If , an event was classified as a left single, if it was a right single, and if it was determined to be a shared event. Inset plots in Fig. 4 show this based discrimination.
@@ IV Noise Correlation
Figure 5 compares this model to the measured noise, showing excellent agreement with only three degrees of freedom (), and supporting the hypothesis that excess sub-threshold events create our observed excess noise. Sub-threshold LEE events which deposit energy instantaneously and locally within the sensor films would produce flat uncorrelated shot noise, while sub-threshold LEE events which generate a shared athermal phonon response would produce correlated shot noise with the observed frequency dependence.
V Discussion
Our observations are consistent with excess events above and below threshold creating backgrounds and shot noise respectively. With excess correlated noise and shared backgrounds, we observe phonon mediation and strong coupling to both channels. Similarly, we observe uncorrelated noise and singles which strongly couple to one channel and are inconsistent with substrate phonon events. If these excess noise terms are indeed LEE shot noise, LEE limits both resolution and backgrounds in our detector. Ultimately, mitigating this noise would allow us to achieve 60 meV sensitivities to phonon events, nearing the level needed to search for single optical phonons generated by dark matter interactions Knapen et al. (2018).
Given the multiple classes of excesses, it seems plausible to attribute these events to multiple sources. For example, aluminum relaxation would naturally explain singles, while shot noise from GHz scale EMI bursts might dominate the uncorrelated noise. Likewise, different effects could dominate phonon backgrounds at different energy scales, e.g., high energy shared LEE might originate from the relaxation of radiation induced defects in the substrate Nordlund et al. (2024), while excess correlated noise could be caused by the absorption of 40 meV photons emitted by the detector circuit boardDu et al. (2022). We leave disentangling these hypothesized contributions to future work studying excess rates and noise over time and their scaling with properties of the detector and surrounding materials.
@@ V Discussion
VI Acknowledgements

### v2

@@ Abstract
We describe observations of low energy excess (LEE) events, background events observed in all light dark matter direct detection calorimeters, and noise in a Transition Edge Sensor based two-channel silicon athermal phonon detector with 375 meV baseline energy resolution. We measure two distinct LEE populations: “shared” multichannel events with a pulse shape consistent with substrate athermal phonon events, and sub-eV events that couple nearly exclusively to a single channel with a significantly faster pulse shape. These “singles” are consistent with events occurring within the aluminum athermal phonon collection fins. Similarly, our measured detector noise is higher than the theoretical expectation. Measured noise can be split into an uncorrelated component, consistent with shot noise from small energy depositions within the athermal phonon sensor itself, and a correlated component, consistent with shot noise from energy depositions within the silicon substrate’s phonon system.
@@
The search for dark matter (DM) has expanded to lower mass candidates, including sub-GeV “light mass” DM Kuflik et al. (2016, 2017); Hochberg et al. (2014, 2015); Hall et al. (2010). Direct detection of light mass DM scattering off nuclei, electrons, or crystal lattices requires extremely low energy thresholds, given the low kinetic energy carried by the DM particles. Cryogenic calorimeters are well suited to attaining such low thresholds, and have recently set limits on sub-GeV DM-nucleon interaction cross sections Abdelhameed et al. (2019); Alkhatib et al. (2021); Angloher et al. (2023a). These calorimeters typically read out athermal phonons from a crystalline substrate using Transition Edge Sensors (TES) Irwin and Hilton (2005) connected to aluminum athermal phonon collection fins (forming structures known as Quasiparticle-trap-assisted Electrothermal-feedback TESs: QETs Irwin et al. (1995)).
@@
Calorimetric DM direct detection experiments and other low threshold calorimeters have observed an excess of events below several hundred eV, with a rate that rises dramatically at low energies Adari et al. (2022); Alkhatib et al. (2021); Abdelhameed et al. (2019). The rate of these low energy excess (LEE) events decreases with time, and can be regenerated by warming up the detector Angloher et al. (2023b). Additionally, the LEE rate varies only weakly with detector material or mass Angloher et al. (2023b), and appears similarly in detectors run above and below groundPyle .
@@
The decrease in LEE rate with time suggests a relaxation mediated process. Mechanical stress relaxation in the detector holding has been shown to create LEE-like events Astrom et al. (2006); Anthony-Petersen et al. (2024); however, a LEE population remains even in detectors held in low stress configurations Anthony-Petersen et al. (2024) (when not discriminating between “singles” and “shared” LEE, as we do here), implying additional relaxation processes are necessary to explain observations.
@@
Stress created by the thermal contraction of sensor films relative to thick detector substrates has been proposed as another LEE source Anthony-Petersen et al. (2024). This stress would be present in all calorimeters observing the LEE, largely independent of the detector material or size.
@@
If relaxation within films is responsible for the LEE, we expect some partitioning of energy between local deposition in the film’s electronic system and phonon energy that leaks into the substrate. For example, athermal phonon bursts within the aluminum QET should break Cooper pairs and locally deposit energy Kozorezov et al. (2000); Martinis (2021). A detector instrumented with multiple individually-read-out athermal phonon sensors would measure this localized energy deposition and therefore be able to discriminate LEE events (with significant local energy absorption) from DM interactions in the bulk substrate (that approximately uniformly excite all phonon sensors). This detector architecture has previously been shown by CRESSTAngloher et al. (2024a) and contemporaneously by our group Romani (2023) to have great potential. Around this time, a version of the film relaxation model briefly proposed in Ref. Anthony-Petersen et al. (2024) was more fully developed and published Romani (2024). This model attempted to explain LEE phonon bursts through the relaxation of thermally stressed aluminum films on the device surface. During the early experimental workRomani (2023), limited control of systematics and understanding of pulse shapes for locally absorbed events constrained our ability to draw strong conclusions about sources of LEE. These limitations have now been remedied.
@@
To test this LEE discrimination concept, a 1 cm-square, 1 mm-thick silicon substrate was instrumented with two channels of 25 tungsten TESs ( 48 mK Tc, nominally 40 nm thick) connected to aluminum athermal phonon collection fins (QETs Irwin et al. (1995), 600 nm thick), covering 1.38 of the device’s surface (see Fig. 1). QETs were electrically connected by partially overlapping their aluminum fins, such that each channel of 25 TESs was read out as one unit. Unfortunately, during manufacturing, a fraction of the TESs were partially etched away, leading to some performance degradation (higher normal resistance and worse phonon collection efficiency). See supplementary material section C and Fig. S4 for further discussion. While performance in both channels was still acceptable, we focus on the left channel, which was less negatively impacted. As in Ref. Anthony-Petersen et al. (2024), this detector was suspended from wire bonds to minimize LEE-type backgrounds associated with detector holding. It was housed inside multiple layers of electromagnetic interference (EMI) and infrared (IR) shielding at the base stage of a dilution refrigerator, and was read out using DC SQUID array amplifiers.
@@
The detector was calibrated with optical photons to characterize its response to events of a known energy (see Figure 2). Pulses of photons from a 405 nm (3.061 eV) room temperature laser were transmitted to the device using a single mode optical fiber terminated with a diffuser, which dispersed photons across the entire instrumented side of the detector. The photon pulses were fast ( 1 s) compared to the electrical ( 10 s) and electrothermal ( 100 s) response times of the TESs. On average photons hit the detector per pulse. We recorded calibration data for 3.5 hours, firing laser pulses at 100 Hz, ultimately recording 1.26 million calibration events. Immediately following the calibration, three hours of background data were acquired, interleaved with periods where the TESs were characterized using IV and dIdV measurements Irwin and Hilton (2005); Watkins (2022).
@@
Both datasets were recorded continuously and triggered offline. For the calibration dataset, we triggered on a recorded logic signal in coincidence with the laser pulses. In the background dataset, events were triggered on the sum of the two detector channels using an optimum filter energy estimator Golwala (2000). See supplementary material section A for further discussion of triggering.
@@
Standard quality cuts were applied to the triggered data to ensure that the detector was operating stably, and to reject periods of high environmental noise or abnormal device performance. These cuts were designed for high passage of randomly triggered events (79.5 %) and for similar passage of high energy events to minimize selection biases. See supplementary material section B for additional discussion. The height of each event was fit with an optimal filter assuming a calibrated phonon pulse shape (see below), and was converted into an energy by applying a factor derived from the channels responsivity Romani et al. (2024) modeled from the measured complex impedance Lindeman et al. (2011); Fink et al. (2020). Using this method, we estimate the amount of energy each event deposits in a given TES channel for an assumed pulse shape (rather than into the detector phonon system as a whole). We use this as the primary quantity we plot in Figs. 2, 3 and 5 as a subclass of events we observe clearly do not couple through the detector phonon system (“singles,” see below). See supplementary material section C for further discussion of energy reconstruction.
@@
In the calibration and background datasets (see Figs. 2 and 3), “shared” events couple roughly equally to both channels (diagonal band, left panels). These events feature a relatively slowly rising pulse (see top right panels) which can be well-modeled as the sum of two exponential phonon pulses convolved with the detector responsivity (). Calibration and background events are identically shaped.
@@
We associate these events with bursts of athermal phonons from the substrate which couple roughly equally to both channels. (Phonons couple somewhat more strongly to the left channel, i.e. have a higher “phonon collection efficiency.” See supplementary material section C for further discussion.) In the background dataset at low energies, we associate these “shared events” with non-sensor film LEE relaxation sources due to the lack of significant localized energy absorption within the channel. At high energies (group of events above 5 eV in the left channel in Fig. 3), the saturated event rate is roughly consistent with the expected rate of high energy events from environmental radioactivity and cosmic rays. See supplementary material section D for additional discussion.
@@
In the calibration, these events are caused by photons absorbed in the substrate, creating quantized (0, 1, 2… photons absorbed) bursts of athermal phonons. We combine the response in both channels using inverse variance weightingRomani et al. (2024), constructing a phonon energy estimator, and plot a combined calibration histogram (see Fig. 2, inset, and material section C for more details on this energy reconstruction). From this histogram, we measure a world leading baseline phonon energy resolution of meV (stat.), favorably comparing to previous detectorsAngloher et al. (2024b); Michielin (2024), both with eV resolution.
@@
Another class of events exhibits a nearly maximally asymmetric channel response (vertical and horizontal bands in Figs. 2 and 3), which we call “single” events due to their strong coupling to a single channel. In the calibration dataset, we attribute this response to events in which one photon hits an aluminum phonon collection fin, i.e. a “direct hit” event. Additional photons may be absorbed in the detector substrate, creating a superposition of “direct hit” and “substrate” events and forming the structure of black bands in Fig. 2. Of the roughly 4.1 single photon events seen in the substrate, we expect 1 will hit a fin in a given channel, with 10 of these photons being absorbedFlaschmann et al. (2023). 239 events fall into the purple box in Fig. 2 (which contains roughly half of the left channel singles), in general agreement with the expected number of events. We attribute the spread in reconstructed energy to instrumental effects (e.g. position dependence within the aluminum, saturation from partial TESs etching), as well as the difference in the apparent distribution of singles energies (see supplementary material section C for further discussion). Given only 0.1 of the QET area is exposed tungsten, we do not expect to observe a significant number of tungsten direct hits.
@@
Pulse shapes for single events are shown in the bottom right panels of Figs. 2 and 3. The fast rise compared to substrate events (though somewhat slower than the modeled TES response to Dirac-delta energy impulses) indicates that the substrate phonon collection dynamics are bypassed. We attribute the slow fall of these events to saturation effects. Quasiparticles created by localized photon absorption would be expected to propagate to only a few of the 25 QETs in a channel, saturating these few TESs at energies significantly below the 5-6 eV of absorbed energy required to saturate all 25 QETs in a channel (see the group of saturated events at high energies in Fig. 3).
@@
In the left singles events in Fig. 4, we see a small average response in the opposite (right) channel, carrying a few percent of the main channel’s energy, in addition to the large average response in the primary (left) channel. The average opposite (right) channel responses for background and calibration events are indistinguishable, and consistent with the calibrated athermal phonon response. We attribute this signal to athermal phonons that leak out of the aluminum fin during downconversion following a singles event. Notably, without rescaling the average background and calibration pulse shapes seen in the opposite (right) channel, these responses are roughly equal in height, suggesting that lower energy background events are more efficient in producing athermal phonons than calibration events, and hinting that background singles may originate deeper within the aluminum fin.
@@
Bursts of high frequency (MHz-GHz) EMI Romani et al. (2024) cannot be the primary source of these background singles, as individual EMI photons would not be sufficiently energetic to leak above the aluminum gap ( ( 90 GHz) ) phonons into the substrate during the downconversion process. We also reject relaxation or other events occurring in the tungsten TESs as the source of background singles, as tungsten relaxation would be expected to produce a far larger pulse of athermal phonons in the substrate due to thinness of the tungsten film ( nm)Romani et al. (2024).
@@
Alternatively, the relaxation of thermally stressed aluminum films could create singles backgrounds. Romani Romani (2024) described a model where relaxing dislocations in an aluminum film impact the aluminum-substrate interface, injecting bursts of athermal phonons into the substrate while depositing minimal energy in the aluminum (i.e. forming shared events). Clearly, observed singles are in the opposite limit: they couple almost exclusively to the aluminum film. Modifying the model in Ref. Romani (2024) to include e.g. damping of dislocations in the bulk film through the emission of above-gap phonons or phonon-emitting interactions with intra-film grain boundaries might better explain singles events.
The background spectra for single and shared events are plotted in Fig. 5. Shared events were triggered on the sum of the two channels using a phonon template, while singles were triggered in the left channel using an averaged singles template. Since a given single or shared event could trigger both single and shared triggers, a statistic considering the pulse shape and amplitude in both channels was used to discriminate event types and avoid double counting. If , an event was classified as a left single, if it was a right single, and if it was determined to be a shared event. Inset plots in Fig. 5 show this based discrimination.
@@
Figure 6 compares this model to the measured noise, showing excellent agreement with only three degrees of freedom (), and supporting the hypothesis that excess sub-threshold events create our observed excess noise. Sub-threshold LEE events that deposit energy instantaneously and locally within the sensor films would produce flat uncorrelated shot noise, while sub-threshold LEE events that generate a shared athermal phonon response would produce correlated shot noise with the observed frequency dependence.
Our observations are consistent with excess events (i.e. LEE) above and below threshold creating backgrounds and shot noise respectively. With excess correlated noise and shared backgrounds, we observe phonon mediation and strong coupling to both channels. Similarly, we observe uncorrelated noise and singles that strongly couple to one channel and are inconsistent with substrate phonon events. If these excess noise terms are indeed LEE shot noise, LEE limits both resolution and backgrounds in our detector. Ultimately, mitigating this noise would allow us to achieve 60 meV sensitivities to phonon events, nearing the level needed to search for single optical phonons generated by dark matter interactions Knapen et al. (2018).
Given the multiple classes of excesses, it seems plausible to attribute these events to multiple sources. For example, aluminum relaxation would naturally explain singles, while shot noise from GHz scale EMI bursts could dominate the uncorrelated noise. Likewise, different effects could dominate phonon backgrounds at different energy scales, e.g., high energy shared LEE might originate from the relaxation of radiation induced defectsNordlund et al. (2024) or “microfractures”Astrom et al. (2006) within the substrate, while excess correlated noise could be caused by the absorption of e.g. 40 meV photons emitted by the detector circuit boardDu et al. (2022). We leave disentangling these hypothesized contributions to future work studying excess rates and noise over time and their scaling with properties of the detector and surrounding materials.
@@
I Supplementary Material
See the supplementary material for details on our optimum filter based trigger scheme, our data quality cuts, the way in which we reconstruct the energy of triggered events, and the saturated events we see in our detector.
II Acknowledgments